# printacmref=false

### v1

@@ Abstract.
Data wrangling—the process of cleaning, transforming, and preparing data for analysis—is a well-known bottleneck in data science workflows. Existing tools either rely on manual scripting, which is error-prone and hard to debug, or automate cleaning through opaque black-box pipelines that offer limited control. We present Buckaroo, a scalable visual data wrangling system that restructures data preparation as a direct manipulation task over visualizations. Buckaroo enables users to explore and repair data anomalies—such as missing values, outliers, and type mismatches—by interacting directly with coordinated data visualizations. The system extensibly supports user-defined error detectors and wranglers, tracks provenance for undo/redo, and generates reproducible scripts for downstream tasks. Buckaroo maintains efficient indexing data structures and differential storage to localize anomaly detection and minimize recomputation. To demonstrate the applicability of our model, Buckaroo is integrated with the Hopara pan-and-zoom engine, which enables multi-layered navigation over large datasets without sacrificing interactivity. Through empirical evaluation and an expert review we show that Buckaroo makes visual data wrangling scalable—bridging the gap between visual inspection and programmable repairs.
@@
Data wrangling involves a wide range of tasks, including parsing, deduplication, missing value imputation, anomaly detection, and type normalization. These tasks are often executed through manual scripts or ad-hoc spreadsheet operations, introducing two significant problems: (1) errors introduced during wrangling are difficult to detect and propagate silently into downstream models, and (2) the lack of interactivity and visibility in existing tools limits user understanding and trust, especially when dealing with complex group-level anomalies.
Bridging the gap between iterative scripting and visual exploration is essential. Data wrangling is rarely a one-shot operation—it is inherently iterative, involving trial-and-error, context-aware corrections, and domain judgment Shastri2023saga. While prior work has proposed automated or script-based approaches to support cleaning operations xubook, these systems often assume a fixed pipeline and offer little visibility into how each transformation affects the dataset, i.e., fixing one error might lead to new errors in the dataset. This limits their usability in real-world settings where errors are heterogeneous, subtle, and contextual.
Conversely, visualization tools excel at surfacing structure vistrails; ruddle2023tasks, guiding attention, and enabling human reasoning, but they are typically detached from the wrangling process itself. That is, they are used post-hoc for validation rather than as integral components of repair. This disconnect leads to workflows where users must constantly switch between error inspection and editing scripts to fix problems in the data, which is laborious and error-prone. Our proposed system, Buckaroo, bridges this gap by reimagining data wrangling as a visual, interactive, and iterative process. It allows users to identify and resolve data anomalies by directly manipulating visual representations of data groups. Buckaroo automatically detects anomalous groups—e.g., those with missing values or outliers—maps them to interactive charts, and offers recommended repairs that can be applied, visualized, and reverted in real time. This tight integration of detection, visualization, and repair enables users to understand the impact of each action across the dataset and supports the inherently exploratory nature of data preparation.
@@
Transforming visualizations into repair interfaces. A key technical insight of Buckaroo is to treat visualizations not merely as passive outputs but as active substrates for user-driven data transformation. By constructing index structures that link anomalies to data groups and anomaly types, Buckaroo enables responsive, bidirectional interactions: users can trigger repairs through visual selections and observe the systemic consequences instantly. Moreover, through user-defined detector and repair functions, the system accommodates domain-specific anomalies while preserving flexibility and reproducibility through automatic code generation.
@@
By enabling users to see, understand, and repair anomalies through a single, unified visual interface, Buckaroo represents a paradigm shift in how practitioners wrangle data. It transforms data wrangling from a brittle and opaque task into an intuitive, transparent, and reproducible process.
@@ 5. Related work
Efforts to improve data quality and make data preparation more accessible have evolved along two trajectories: (1) Data cleaning research xubook has produced numerous techniques for detecting and repairing errors; however, little attention was given to user-friendly interfaces for data cleaning. (2) Visualization and visual analytics systems provide rich, interactive substrates for examining data, yet they are predominantly employed in a post-hoc, observational capacity rather than as active components of the data cleaning process. Buckaroo operates at the intersection of these two lines of work, integrating the algorithmic foundations of data cleaning with direct visual manipulation to enable users to visually inspect, adjust, and steer repairs as part of an exploratory workflow.
@@ 5.2. Visualization
A range of systems have investigated how to visually support the data wrangling process potterwheel. For example, xiong2022revealing; xiong2022visualizing explore techniques for visualizing wrangling scripts. Buckaroo takes the inverse approach by embedding wrangling actions directly within visualizations. kasica2020table developed a framework describing the wrangling operations available to multi-table data, specifically in data journalism contexts, which was subsequently used by xiong2022revealing. Similarly, Wrangler kandel2011wrangler—later commercialized as Trifacta—synthesizes transformation scripts based on user interactions, a concept recently extended with natural language prompting in Dango chen2025dango. Profiler kandel2012profiler subsequently developed statistical displays of collections of errors across a dataset, a design which we draw on in Buckaroo. shrestha2021unravel develop a method for visualizing and manipulating data frame wrangling operations governed by fluent interfaces like Pandas. zhu2025visegpt explore how visualization can be used to support understanding automatically generated data wrangling scripts through a bespoke Gantt chart. Scorpion wu13scorpian draws on user-identified outliers to synthesize predicates that explain outliers. Unlike these systems, which rely on programming-by-demonstration, Buckaroo emphasizes real-time, direct manipulation of visual elements to guide repairs.
These systems in essence are a specialized form of visualization recommendation system (for which zeng2024systematic provide an overview), in which the gestural input to that recommendation system that is used to drive the recommendation is the wrangling process itself. Our work draws on this tradition with a more specific focus. Other systems in this area bear a connection with ours. For instance, Data formulator wang2023data; wang2025data combines AI-based natural language guidance with a closely related visualization-by-demonstration wang2021falx technique, which observes that preparation and mapping are interwoven components of the presentation process. To that end, Lux lee2021lux weaves visualization recommendation into notebooks, in a manner which centers a predefined collection of analysis tasks. We invert this design, by placing wrangling inside of a visualization environment, rather than placing visualizations in a data programming context.
@@
Lastly, our work draws on the long history of visual analytics systems connected with data management systems. These include a wide array of different tools and systems battle2020structured; wu2014case. Of particular relevance are GUI-based visual analytics systems, such as Polaris stolte2002polaris (subsequently commercialized as Tableau). Our work draws on the patterns and traditions of these systems, but focuses on the particular task of data wrangling.
6. Preliminary results and next
steps

### v2

@@ Abstract.
Data wrangling—the process of cleaning, transforming, and preparing data for analysis—is a well-known bottleneck in data science workflows. A wide range of data wrangling techniques have been proposed to mitigate this challenge. Of particular interest are visual data wrangling tools, in which users prepare data via graphical interactions (such as with visualizations) rather than requiring them to write scripts. We develop a visual data wrangling system, Buckaroo that expands upon this paradigm by enabling the automatic discovery of interesting groups (e.g., Salary values for Country=“Buthan”) and identification of anomalies (e.g., missing values, outliers, and type mismatches) both within and across these groups. Crucially, this allows users to reason about how repairs applied to one group affect other groups in the dataset.
A central challenge in visual data wrangling is scalability. Rendering entire datasets is often infeasible, yet showing only a small sample risks hiding rare but critical errors across groups. We address these challenges through carefully designed sampling strategies that prioritize errors, as well as novel aggregation techniques that support pan-and-zoom interactions over large datasets. Buckaroo maintains efficient indexing data structures and differential storage to localize anomaly detection and minimize recomputation. We demonstrate the applicability of our approach via an integration with the Hopara pan-and-zoom engine (enabling multi-layered navigation over large datasets without sacrificing interactivity). Finally, we explore our system’s usability (via an expert review) and its scalability, finding that this design seems well matched with the challenges of this domain. [^1] This version revises the paper to improve its positioning with respect to prior work in response to feedback received after publication. Earlier versions remain available in the arXiv version history.
@@
Data wrangling involves a wide range of tasks, including parsing, deduplication, missing value imputation, anomaly detection, and type normalization. Further complicating these challenges is that fixing one data anomaly can lead to other anomalies. Several approaches have been proposed to mitigate error detection and correction readytodeploy. In particular, visual data wrangling techniques kandel2011wrangler; kandel2012profiler enable users to identify and correct data anomalies through graphical interactions such as visualizations. We present Buckaroo, which extends this paradigm by introducing three key capabilities for visual data wrangling: (1) Group-level anomaly detection, where Buckaroo automatically generates interesting data groups (e.g., salary values for Country = “Bhutan”) and detects anomalies both within and across groups (e.g., a value may be an outlier in one group but not in another); (2) a pan-and-zoom interface that allows users to focus on regions of interest within a visualization; and (3) anomaly-centric sampling and aggregation, which provides multiple sampling and aggregation strategies to support scalable visual data wrangling on large datasets.
Data wrangling for anomaly detection is rarely a one-shot operation—particularly when anomalies are subgroup-dependent. An error that appears anomalous in one subset of the data may be expected or even valid in another, necessitating iterative, context-aware inspection and correction informed by domain knowledge. In this work, we propose to (1) identify semantically meaningful data groups (e.g., Education for Country = “Bhutan”), (2) detect anomalies within and across these groups, and (3) support their repair through scalable visual interfaces.
It is well known that visualization tools excel at surfacing structure in data vistrails; ruddle2023tasks and supporting data wrangling tasks kandel2011wrangler; kandel2012profiler. We build on this line of work by enabling data wrangling for subgroup anomaly detection and correction through interactions with interactive charts. Our proposed system, Buckaroo, frames subgroup data wrangling as a visual, interactive, and iterative process. It allows users to identify and resolve group-level data anomalies by directly manipulating visual representations of data groups. Buckaroo automatically detects anomalous groups—e.g., those with missing values or outliers—maps them to interactive charts, and offers recommended repairs that can be applied, visualized, and reverted in real time. This tight integration of detection, visualization, and repair enables users to understand the impact of each action across the dataset and supports the inherently exploratory nature of data preparation.
@@
Transforming visualizations into repair interfaces. A key technical insight of Buckaroo is to treat visualizations as active substrates for user-driven group data transformation. By constructing index structures that link anomalies to data groups and anomaly types, Buckaroo enables responsive, bidirectional interactions: users can trigger repairs through visual selections and observe the systemic consequences instantly. Moreover, through user-defined detector and repair functions, the system accommodates domain-specific anomalies while preserving flexibility and reproducibility through automatic code generation.
@@
By enabling users to see, understand, and repair subgroup anomalies through a single, unified visual interface, Buckaroo represents a paradigm shift in how practitioners wrangle data. It transforms data wrangling from a brittle and opaque task into an intuitive, transparent, and reproducible process.
@@ 5. Related work
Efforts to improve data quality and make data preparation more accessible have evolved along two trajectories: (1) Data cleaning research xubook has produced numerous techniques for detecting and repairing errors; however, little attention was given to user-friendly interfaces for data cleaning. (2) Visual data wrangling techniques provide rich, interactive substrates for examining data and wrangling data visually. Buckaroo operates at the intersection of these two lines of work, integrating the algorithmic foundations of data cleaning with direct visual manipulation to enable users to visually and scalably inspect, adjust, and steer repairs for subgroup analysis as part of an exploratory workflow.
@@ 5.2. Visualization
We build on prior work in interactive data wrangling potterwheel and self-service data preparation Hellerstein2018SelfServiceDP. Wrangler kandel2011wrangler; wrangler_cidr—later commercialized as Trifacta—introduced a direct-manipulation paradigm in which transformation scripts are synthesized from user interactions with graphical interfaces, including visualizations. Buckaroo follows this paradigm but shifts the focus to subgroup anomaly detection and correction. In addition, Buckaroo introduces error-centric sampling and aggregation strategies to support scalable analysis. While Trifacta also provides sampling capabilities alteryx_sampling_overview_2025, available documentation suggests that these strategies are not explicitly designed around error detection. An important follow-up work is Profiler kandel2012profiler, which subsequently developed statistical displays of collections of errors across a dataset, a design which we draw on in Buckaroo. Buckaroo is complementary to Profiler in that it provides support for subgroup anomaly analysis.
Another recent work is Dango chen2025dango which extends visual data wrangling with natural language prompting. xiong2022revealing; xiong2022visualizing explore techniques for visualizing wrangling scripts. Buckaroo takes the inverse approach by embedding wrangling actions directly within visualizations. kasica2020table developed a framework describing the wrangling operations available to multi-table data, specifically in data journalism contexts, which was subsequently used by xiong2022revealing. shrestha2021unravel develop a method for visualizing and manipulating data frame wrangling operations governed by fluent interfaces like Pandas. zhu2025visegpt explore how visualization can be used to support understanding automatically generated data wrangling scripts through a bespoke Gantt chart. Scorpion wu13scorpian draws on user-identified outliers to synthesize predicates that explain outliers. Buckaroo falls into the lineage of these systems, drawing on many of the usage patterns demonstrated in those systems. Our work is differentiated from this previous ones through our focus on subgroup analysis.
Many visual data wrangling systems can be seen as a specialized form of visualization recommendation system (for which zeng2024systematic provide an overview), in which the gestural input to that recommendation system that is used to drive the recommendation is the wrangling process itself. Our work draws on this tradition with a more specific focus. For instance, Data formulator wang2023data; wang2025data combines AI-based natural language guidance with a closely related visualization-by-demonstration wang2021falx technique, which observes that preparation and mapping are interwoven components of the presentation process. Lux lee2021lux weaves visualization recommendation into notebooks, in a manner which centers a predefined collection of analysis tasks. We invert this design, by placing wrangling inside of a visualization environment, rather than placing visualizations in a data programming context.
@@
Lastly, our work draws on the long history of visual analytics systems connected with data management systems. These include a wide array of different tools and systems battle2020structured; wu2014case. Of particular relevance are GUI-based visual analytics systems, such as Polaris stolte2002polaris (subsequently commercialized as Tableau). Our work draws on the patterns and traditions of these systems, but focuses on the subgroup analysis within visual data wrangling.
6. Preliminary results & next steps
@@ 6.3. Next Steps and Concluding Remarks
Acknowledgments
We sincerely thank the anonymous reviewers, Michael Stonebraker, and Joseph Hellerstein for their insightful and constructive comments, which greatly improved both the project and the paper.